OpenStudy (anonymous):

Given the equation 3x - 3 = 6x + 1, select the reasoning that correctly solves for x. Add 3, subtract 3x, then divide by -3 Add 3, subtract 6x, then divide by 3 Subtract 1, subtract 3x, then divide by 3 Subtract 1, subtract 6x, then divide by -3 Which one Is it? Also I'd like an explanation

OpenStudy (texaschic101):

3x - 3 = 6x + 1 --- add 3 3x = 6x + 1 + 3 3x = 6x + 4 --- subtract 6x -6x + 3x = 4 -3x = 4 -- divide by -3 x = -4/3